OpenStudy (anonymous):

In the equation y = 1/4x + 8 , what is y when x is 20 ?

OpenStudy (anonymous):

y = 1/4x + 8 x=20 y = 1/4(20) + 8 y=13

OpenStudy (anonymous):

Good luck:D

OpenStudy (anonymous):

I love you!

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Which of the following equations is graphed as a straight line? y = 3x y= 4/x y = 7x to the 2nd power or xy= 8

OpenStudy (anonymous):

y = 3x has a straight line, i just finished graphing